OpenStudy (anonymous):

I need help solving a couple of Beginning Algebra equations. -1/3 = x - 3/5

OpenStudy (lollylau):

you wanna get rid of any fractions by multiplying all the terms by their denominators. so the whole equation becomes: -5=15x-9. -15x=-4 15x=4 x=4/15

OpenStudy (anonymous):

So -1/3 . - 3/5 = 3/15 The fractions is what gets me. But thank you.

OpenStudy (anonymous):

this is so you can cancel the -3 / 5

OpenStudy (anonymous):

Ok so you multipy on both sides the -1/3

OpenStudy (anonymous):

multiply it with a - 5/ 3

OpenStudy (anonymous):

you have -1/3 = x - 3/5 it would be better to multiply it in 1 go and simplify

OpenStudy (anonymous):

multiply with reciprocal of -3/5 it would be easier you could do -1/3 reciprocal but it would be more work

OpenStudy (anonymous):

So either side is ok? I didn't know that. Thank you!

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

any other questions?

OpenStudy (anonymous):

Yes I have another equation 3x = 2x + 7

OpenStudy (anonymous):

since if you multiply for example a equation by -1 you have to do so to both sides you know that right?

OpenStudy (anonymous):

No Im so lost right now.

OpenStudy (anonymous):

o okies you know what this sign = means right?

OpenStudy (anonymous):

equals yes lol

OpenStudy (anonymous):

it means is equal to

OpenStudy (anonymous):

the equation 3x=2x+7 it reads 3x is equal to 2x plus 7

OpenStudy (anonymous):

so do you kinda understand what this is know

OpenStudy (anonymous):

now*

OpenStudy (anonymous):

how it works

OpenStudy (anonymous):

Yes I know we have to combine like terms but when I did it I came up with the wrong answer

OpenStudy (anonymous):

I came up with 1x

OpenStudy (anonymous):

not like terms but equal amounts just think of it like this you got a 2 ml cups of water A and B they are supposed to be equal if you add 4 pints to cup A Will they still be equal? How will you make them equal? Add 4 pints to cup B right? what is the symbol of adding? its a plus sign + 4 pints

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it more understandable?

OpenStudy (anonymous):

a little so do i add the 2x to 3x or am I still off base?

OpenStudy (anonymous):

2x = 3x + 7 what would be the easiest method try to make it simpler for you self yes , correct make the like terms dissapear

OpenStudy (anonymous):

I did it like this: 3x = 2x + 7 -2x -2x

OpenStudy (anonymous):

I know the answer is 7 from looking in the back of the book just can't figure out how to get the answer to 7

OpenStudy (anonymous):

additive inverse of 2x is negative 2x do this on both sides

OpenStudy (anonymous):

yes I did that

OpenStudy (anonymous):

x cannot be negative

OpenStudy (anonymous):

how would you get a positive x?

OpenStudy (anonymous):

I"m not sure but I think the +3x -2x should egual a positive but that is 5x

OpenStudy (anonymous):

umm you sure the problem was 2x = 3x + 7

OpenStudy (anonymous):

not - 7

OpenStudy (anonymous):

no the problem is 3x =2x + 7

OpenStudy (anonymous):

answer is 7 but don't see how to get that answer.

OpenStudy (anonymous):

okies here is what i get 3x = 2x +7 (-2 x) 3x = 2x + 7 + (-2x) additive inverse of 2x

OpenStudy (anonymous):

can you figure the rest :)

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

anything else?